OpenStudy (anonymous):

"1/R effective = 1/12ohms + 1/15ohms" And just so you know, it's like the ohms symbol is in the denominator with the numbers, not to the side of it. The answer is 6.67 ohms, but I don't understand how they got there. Could you please solve/explain this?

OpenStudy (jamesj):

If \[ \frac{1}{R} = \frac{1}{12} + \frac{1}{15} \] you need to solve for R. Can you do that?

OpenStudy (anonymous):

first you need a common denominator \[\frac{1}{R_{eff}} =( \frac{1}{12} \times \frac{5}{5} )+(\frac{1}{15} \times \frac{4}{4})\]\[\frac{1}{R_{eff}}=\frac{5}{60}+ \frac{4}{60}\]\[=\frac{4+5}{60} = \frac{9}{60}\]so \[\frac{1}{R_{eff}}= \frac{9}{60}\]\[\rightarrow R_{eff} = \frac{60}{9}\]

OpenStudy (anonymous):

^^ thanks so much! that was exactly what i needed! :D

OpenStudy (anonymous):

=D your welcome